OpenStudy (edaiiiii):

about how many pounds are there in 4.5 kilograms?

OpenStudy (astrophysics):

\[ 1 kg = 2.2 lbs\]

OpenStudy (edaiiiii):

ok

OpenStudy (astrophysics):

Use unit conversions, \[\left( 4.5 kg \right)\left( \frac{ 2.2 lbs }{ 1 kg } \right)\]

OpenStudy (edaiiiii):

so what i gotta do

OpenStudy (astrophysics):

Multiply

OpenStudy (edaiiiii):

what am i multiplying

OpenStudy (liziekay):

distribute

OpenStudy (liziekay):